OpenStudy (anonymous):

For questions 5-7, find the solution to the system of equations by using either graphing or substitution. 6. y=2x-1 and y=x+3 A. (4, 7) B. (7, 4) C. (-7, -4) D. Infinite Solutions

OpenStudy (jdoe0001):

so you have $$ 2x-1=y\\ \color{blue}{x+3=y}\\ --------- $$

OpenStudy (anonymous):

I believe the answer is A.... Is that correct?

OpenStudy (jdoe0001):

well, what would you need in the "blue" or bottom equation, to have it substract from the top and get "x" by itself?

OpenStudy (anonymous):

Idk... I'm not very knowledgeable about algebra.

OpenStudy (jdoe0001):

well, I assume you were given this because it has already being covered by the textbook, and you're supposed to know it

OpenStudy (jdoe0001):

otherwise, you shouldn't be doing it

OpenStudy (anonymous):

Mhm. And I have been trying to figure out the solution my textbooks solution is just a little complicated for me to understand..

OpenStudy (jdoe0001):

lemme retype the equations some

OpenStudy (jdoe0001):

$$ 2x-1=y\\ \color{blue}{x\ +\ 3=y}\\ ---------\\ \color{blue}{x\ +\ ?=\ ?}\\ $$

OpenStudy (jdoe0001):

you'd want the 2nd equation, the bottom one, to "turn into" something that when added to the top one, gives only one variable, and cancels out the other

OpenStudy (jdoe0001):

so you may end up with something like 3x+5=0, or 0+8=y, or so

OpenStudy (jdoe0001):

for this instance, so you can see it, try $$ 2x-1=y\\ \color{blue}{x\ +\ 3=y} \ \ \times (-1)\\ ---------\\ \ \ \ 2x-1=y\\ \color{blue}{-x-3=-y}\\ ---------\\ $$ ? - ? = ?

OpenStudy (jdoe0001):

add them both "vertically", what do you get?